AprilandNate:

Hamish and Harry work as plumbers. Harry earns a dollar more than 5/4 the amount Hamish earns per hour. The amount Harry earns per hour is $2 less than 7/5 the amount Hamish earns per hour. How much does each of them earn per hour?

jhonyy9:

i think may be these wrote Harry earns 1+ (5/4)x the amount Hamish earns per hour and Harry earny per hour (7/5)x -2 the amount Hamish earny per hour the amount what harry earns per hour note with y and in this way you can writing y = 1 +(5/4)x and y = (7/5)x -2 using these knowladges you can writing 1+(5/4)x = (7/5)x -2 from this equation you need calcule the value of x so what will be the amount what Hamish earns per hour @Tranquility how you think please these above wrote may be right,correct ? thank you

Tranquility:

I concur with Jhonny. He deduced the correct equations from the statements provided. Once you solve for x (how much Hamish earns per hour), plug it back into either equation and solve for y, which is how much Harry earns per hour

jhonyy9:

ty @Tranquility